OpenStudy (anonymous):

Rewrite 9/10 and then 23/25 so they have a common denominator. Then use > or < or = . Can someone please help. I really need help with this one :(

jimthompson5910 (jim_thompson5910):

the LCM of 10 and 25 is 50 notice how if we lay out the multiples of 10 and 25, we get 10: 10, 20, 30, 40, 50, 60, 70, etc 25: 25, 50, 75, 100, etc

jimthompson5910 (jim_thompson5910):

do you see how I got 50?

OpenStudy (hockeychick23):

(9/10)*(5/5)= 45/50 (23/25)*(2/2)= 46/50 9/10<23/25

OpenStudy (anonymous):

Thank you hockeychick23